OpenStudy (anonymous):

why d/dx tanx= sinx/cosx and d/dx tanx= sec^2x why both different

OpenStudy (anonymous):

I'll show you trigonometric derivatives

OpenStudy (anonymous):

only tanx= sinx/cosx where D/DX tanx=sec^2x

OpenStudy (anonymous):

ryt??

OpenStudy (anonymous):

noo d/dx sinx is not tanx sinx/cosx=tanx :)

OpenStudy (displayerror):

Remember these two derivatives: \[\frac{d}{dx}\sin (x) = \cos(x)\] \[\frac{d}{dx}\cos (x) = -\sin (x)\] and these two trig identities: \[ \sec (x) = \frac{1}{\cos (x)} \] \[ \sin^2 (x) + \cos^2 (x) = 1\] Like you said, \[\tan (x) = \frac{\sin (x)}{\cos (x)} \] To take the derivative of \( \tan (x) \), you have to use the quotient rule: \[\frac{d}{dx} \frac{f(x)}{g(x)} = \frac{g(x) f'(x) - f(x) g'(x)}{\left(g(x)\right)^2}\] Applying the quotient rule to \(\tan (x)\): \[\frac{d}{dx} \frac{\sin (x)}{\cos (x)} = \frac{\cos (x) \cos (x) - \sin (x) \left(- \sin (x) \right) }{\cos^2 (x) }\] This simplifies to \[\frac{d}{dx} \frac{\sin (x)}{\cos (x)} = \frac{\cos^2 (x) + \sin^2 (x) }{\cos^2 (x)}\] Further simplification with the identities mentioned above will show that \[\frac{d}{dx} \tan (x) = \sec^2 (x)\]